Session: Nov 2004 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C. Bortolotti at Thu Oct 28 04:04:50 2004) Success :

Onsala (Michael Lindqvist at Thu Oct 28 07:12:13 2004) Minor failures :
Problems with the noisecal (i.e, bad tsys-values) during some scans. It did not fire. Can be seen if one compare the tpi/tpical-values where the tsys-valus are bad. Otherwise no known problems.

Noto (P.Cassaro at Thu Oct 28 11:09:56 2004) Minor failures : TSYS
The BBC08 gave crazy values of tsys for the whole experiment.

Effelsberg (D.Graham at Fri Oct 29 08:23:16 2004) Minor failures : HIGH
heavy rain after midnight

Jodrell_Bank (Alastair Gunn at Mon Nov 1 00:57:23 2004) Minor failures : WIND
The Lovell Telescope was used. Seemed to lose contact with Mk5 at about 16:36 UT. Back on for scan 16 - few scans lost. Telescope parked due to high winds from 18:55 UT to 01:06 UT. No other known problems.

Westerbork (Tony Foley at Mon Nov 1 11:09:59 2004) Success :
Wb used in phased array mode. Minor problem with the VSN of the pack being wrong (VSN says OSOD-002 but label is MED-0004). All telescopes used